OpenStudy (anonymous):

Regan is trying to find the equation of a quadratic that has a focus of (−2, 5) and a directrix of y = 13. Describe to Regan your preferred method for deriving the equation. Make sure you use Regan's situation as a model to help her understand.

OpenStudy (anonymous):

please help.

OpenStudy (anonymous):

@DanJS

Miracrown (miracrown):

The graph of the quadratic is a parabola. The directrix is horizontal, so the parabola is vertical. The focus lies below the directrix, so the parabola opens downwards. General equation of a down-opening parabola: y = a(x - h)² + k with a < 0 vertex (h, k) focal length p = | 1/(4a) | focus (h, k-p) directrix y = k+p Plug your data into the equations and solve for a.